OpenStudy (kendricklamar2014):

Explain how to write a function rule from the table below. Then write a function rule. X: 2, 4, 6 Y: 1, 0, -1

OpenStudy (amistre64):

i would do newtons method; setup an equation has roots in X, to solve for Y that way the X parts zero out the coefficients 1 by 1 and you end up with a solution

OpenStudy (kendricklamar2014):

Can u help me, i dont know how to do that

OpenStudy (anonymous):

I can try one sec

OpenStudy (anonymous):

um well first of do u know the equation or tht wht ur trying to find?

OpenStudy (kendricklamar2014):

yah, X: 2, 4, 6 Y: 1, 0, -1

OpenStudy (anonymous):

One second, I think I know

OpenStudy (amistre64):

newtons way: x-2=0, when x=2 x-4 = 0 when x=4 x-6 = 0 when x=6, use these to zero out the terms y = a + b(x-2) + c(x-2)(x-4) + ... well the rest are zeros regardless now let y be what it is when x is a value and solve for the a b c parts

OpenStudy (amistre64):

other than that, you can try to fit it to some quadratic with a system of 3 equations in 3 unknowns

OpenStudy (kendricklamar2014):

So: Y-1=0 Y-0=0 Y--1=0

OpenStudy (anonymous):

x is the input and y is the Dependant variable

OpenStudy (anonymous):

to write a function, you need to the gradient and the y intercept

OpenStudy (anonymous):

you need the gradient** Gradient, m= rise/run= -1/4

OpenStudy (anonymous):

y intercept 1= -(1/4)* 2 + c

OpenStudy (amistre64):

y = a + b(x-2) + c(x-2)(x-4) when x=2, y=-1 -1 = a + b(2-2) + c(2-2)(x-4) 0 0 -1 = a therefore: y = -1 + b(x-2) + c(x-2)(x-4) ; let x=4, y=0 0 = -1 + b(4-2) + c(x-2)(4-4) 0 0 = -1 +2b ; b = 1/2 etc

OpenStudy (anonymous):

1= -(1/2) + c

OpenStudy (anonymous):

c = 1 +1/2 = 3/2 Therefore y= -1(1/4)x + 3/2

OpenStudy (anonymous):

I've made a mistake the gradient is -1/2 and thus the y-intercept is 2

OpenStudy (amistre64):

lol, i got my -1 and 1 mixed around ... but Jess idea works out too

OpenStudy (anonymous):

making our function y= -(1/2)x +2 Is that clear?

OpenStudy (kendricklamar2014):

SO the function rule is 1= -(1/2) + c

OpenStudy (anonymous):

The function would be y= -(1/2)x +2

OpenStudy (amistre64):

y = 1 - (x-2)/2 + c(x-2)(x-4), let x=6 and y=-1, solve for c -1 = 1 - (6-2)/2 + c(6-2)(6-4) -2 = -2 + c(4)(2) 0 = 8c, c=0 therefore: y = 1 - (x-2)/2 = 1 - x/2 +1 = -x/2 + 2

OpenStudy (anonymous):

exactly

OpenStudy (kendricklamar2014):

oh ok

OpenStudy (kendricklamar2014):

and that ^ is how u right a function rule for the table?

OpenStudy (anonymous):

We're done now, I hope I've helped :)

OpenStudy (anonymous):

lol I sat back and watched them wrk it out and they got it right man/girl

OpenStudy (kendricklamar2014):

which one is how to write a function rule?

OpenStudy (kendricklamar2014):

@Jesstho.-.

OpenStudy (anonymous):

Scroll up, I've explained everything :)

OpenStudy (kendricklamar2014):

thanks :D

OpenStudy (anonymous):

No problem :)